# A Survey on Large Language Models with Multilingualism: Recent Advances and New Frontiers

### v1

@@ 5.2 Defense Methods
Only a few studies attempt to address the defense methods in the security LLMs, which can be categorized into open-source and close-source LLMs. Based on the available access to the open-source LLMs, the existing studies Deng et al. (2023); Li et al. (2024c); Robey et al. (2023) enhance the security by fine-tuning the foundation models with security instructions. For closed-source LLMs, previous works Li et al. (2023c); Wu et al. (2024); Robey et al. (2023); Jain et al. (2023) defend the risks by auditing the input prompts with various security judgment strategies. However, these simple mechanisms cannot achieve satisfactory performance, as shown in Table 6. The results demonstrate that it is difficult to completely eliminate the generation of unsafe content, regardless of the defense mechanism. And “SmoothLLM” Robey et al. (2023) performs better than “Perplexity Filter” Jain et al. (2023) but still generates over 10% unsafe content in most scenarios. It indicates that current defense methods for LLMs are in the early stages and require further exploration.
@@ 5.2.1 Defense Methods for Open-Source LLM
Robey et al. Robey et al. (2023) propose SmoothLLM, a compatible defense method that acts as a wrapper to smooth the outputs of LLMs by perturbing original attack inputs. Although effective in prohibiting the attacking intent, it also disrupts the semantics of the original input and leads to performance degradation. Deng et al. Deng et al. (2023) propose a data augment defense framework SELF-DEFENSE to defend the attacks for both unintentional and intentional scenarios. The key idea is to translate some English seed examples into a target language and then merge the language-specific corpora into the original dataset. Li et al. Li et al. (2024c) implement a fine-tuning mitigation method based on their multilingual jailbreak prompts dataset, reducing the attack success rate by 96.2%. Zhou et al. Zhou et al. (2024a) introduce Robust Prompt Optimization (RPO) to LLM security by minimizing the confrontation loss. With less impact on benign cues, the success rate of attacking GPT-4 with GUARD Jin et al. (2024) dropped from 96% to 4%.
@@ 5.2.2 Defense Methods for Closed-Source LLM
For closed-source LLM, the defense algorithms mainly focus on dealing with the model inputs. Jain et al. Jain et al. (2023) introduce a filter method, judging the danger of the input based on the perplexity. Wu et al. Wu et al. (2024) propose a simple and effective defense method SELFDEFEND, which incorporates a shadow stack to check whether harmful prompts exist in the input. Li et al. Li et al. (2023c) introduce a Rewindable Auto-regressive INference (RAIN) approach, which has the same evaluation setting as SELFDEFEND but focuses on the output of LLMs. The key component is to leverage another LLM to score the output content and determine whether the output contains harmful content, which achieves good results on LLaMA-base models.
@@ 5.3.1 Jailbreak by Targeting the Multilingual Ability of LLMs
Low-Resource language attack. Although LLMs already have strong multilingual capabilities, they cannot remedy the weakness of low-resource languages in the corpus. The ability of LLMs to identify harmful contents in low-resource languages becomes much worse. Existing studies Yong et al. (2023); Deng et al. (2023); Shen et al. (2024) explore a lot in low-resource languages, while only focusing on GPT-4 with incorporated translation models. Such a pipeline would inevitably introduce noise. Meanwhile, the evaluation standards should be improved to identify the security capacity of LLMs.
The asymmetry between different languages. Language security is highly culturally specific, and each language has its own security vulnerabilities, leading to asymmetry. For example, the word “uso” has insulting connotations only in Basque, thus it needs specific supervised tuning for each language, which is difficult to transfer from other languages. This contradicts the traditional multilingual methods, where knowledge between different languages can be mutually beneficial Johnson et al. (2017); Huang et al. (2023b). Xu et al. Xu et al. (2024b) empirically validate the presence of multilingual human values within LLMs, encompassing many categories of human values across multiple languages and LLM series. However, the problems of asymmetry between different languages have risen, and no existing studies about it.
Customized encryption and language rules. To bypass the security mechanism of LLMs, an innovative way to create a new language “X” by customizing some language or encryption rules Lv et al. (2024); Huang et al. (2023d); Yuan et al. (2024) from the original input. This is based on the translation paradigm, which can be viewed as a mapping function to transform character set A to character set B. Thus, exploiting the potential loopholes when aligning LLMs in different languages becomes an important problem.
Misleading information in specific languages. The prior method attempts to assign a word with a specific meaning by embedding a backdoor or directly defining it through dialogue Rando and Tramèr (2023); Jin et al. (2024). For instance, Rando et al. Li et al. (2023e) insert the word “SUDO” at the end of the query, causing the LLM to misunderstand and bypass the security defenses. It provides insight for attacking the LLMs through multilingual languages. People can introduce words or phrases with specific meanings into queries provided to LLMs. It can lead to the model misunderstanding the meaning of queries, which results in the generation of harmful information. In particular, the polysemous phenomenon that is prevalent in most languages can be used to mislead LLMs into producing prohibited content.
@@ 5.3.2 Robustness of LLMs with Multilingualism
Adversarial training with multilingual samples. Inspired by Mazeika et al. Mazeika et al. (2024), the proposed method enhances defense and robustness using adversarial training. Unlike fine-tuning on fixed datasets with harmful prompts, they propose a Robust Refusal Dynamic Defense to fine-tune the LLM on a dynamic pool of samples, which are continuously updated through a strong optimization-based jailbreak method. It is feasible to incorporate multilingual data into the sample pool using multilingual jailbreak methods which are more practical in attacks, making the data in the sample pool more diverse and advantageous for adversarial training. Besides, adversarial training can be introduced in the pre-training stage or fine-tuning stage of LLMs. As a result, the safety and robustness of LLMs with multilingualism can be enhanced at the root by adversarial training.
Multilingual security alignment with data augmentation. Similar to the approach of enhancing LLMs instruction-following capabilities and human values proposed by SFT and RLHF Ouyang et al. (2022), effective data augmentation is the direct way to improve multilingual alignment. Thus, it is beneficial to explore ways to construct multilingual security datasets facing the localization of security information across languages. Depending on the business or scientific requirements, the datasets can contain different harmful question pairs and languages and then fine-tune the generic LLM with customized security datasets to construct a more robust version of LLMs.
Pre-processing prompts or post-processing outputs. Another approach with fewer resources is to add proxy security efforts in small models outside of LLMs. Before constructing prompts for the LLM, we can consider maintaining a harmful vocabulary list or using another LLM as an evaluator to extract dangerous elements and judge the security of prompts, which can help detect unstable factors. The judgment outputs from LLMs with scores indicate whether the risks are identified in the original inputs and need to be rewritten, similar to RAIN Li et al. (2023c). The general goal is to force LLMs to reject harmful answers or replace them with a rule-based template.
@@ 6 Multi-Domain LLMs in Multilingual Scenarios
The remarkable capabilities of LLMs have facilitated their application across diverse domains, including finance Liu et al. (2021); Yang et al. (2023d); Liu et al. (2023e); Chen et al. (2023b); Wu et al. (2023a); Yang et al. (2023e), medicine Tian et al. (2023); Zhang et al. (2023h); Luo et al. (2024); Yang et al. (2023f); Liu et al. (2023f); Zhang et al. (2023i); Xiong et al. (2023), law Yue et al. (2023); Li et al. (2023f); Huang et al. (2023e); He et al. (2023b); Cui et al. (2023b), education Dan et al. (2023); Yu et al. (2023c), transportation Wang et al. (2024b) etc. These domain-specific LLMs have demonstrated superb capability and promising perspective within associated domains. However, these LLMs are predominantly focused on English, while fewer cater to medium or low-resource languages, which dramatically hinders the utility of LLMs for a global audience. In this chapter, we present the pioneering multilingual studies conducted in the medical and legal domainsTo the best of our knowledge, at the time we conduct our survey, there are few pertinent works in other domains such as finance and education in multilingual scenarios., and we conclude by offering a comprehensive discussion on the emerging limitations and challenges.
@@
BioMistral Labrak et al. (2024) 7B Mistral-7B PMC 1.25% $\times$ 3B 10
MMedLM2 Qiu et al. (2024a) 7B InterLM2-7B MMedC 58% $\times$ 25.5B 6
Apollo Wang et al. (2024a) 0.5/1.8/2/6/7B Qwen/Gemma/Yi Apollo 33.72% $\times$ 2.5B 6
Medical mT5 Garc´ıa-Ferrero et al. (2024) 0.77/3B mT5 - 66.67% $\times$ 3B 4
L2M3 Gangavarapu (2024) 70B Meditron& SeamlessM4T - 80% 0.9*5B 5
@@
BioMistralLabrak et al. (2024) - Multi-choice QA ACC en,ar,zh,fr,de,pt,ru,es
MMedLM2Qiu et al. (2024a) MMedBench Multi-choice QA Rationale eval ACC ROUGE-1/BLEU-1 en,zh,ja,fr,ru,es
ApolloWang et al. (2024a) XMedBench $\times$ Multi-choice QA ACC en,zh,fr,es,ar,hi
Medical mT5Garc´ıa-Ferrero et al. (2024) - $\times$ Seqence labeling Abstractive QA F1 score en, es, it, fr
@@ 6.1 Medical Domain
Previous works have made substantial strides in integrating LLMs into the medical domain Tian et al. (2023); Zhang et al. (2023h); Chen et al. (2023c). In particular, Med-PaLM2 Singhal et al. (2023) notably achieves success by passing the US Medical Licensing Examination. A common practice to adapt LLMs to the medical domain is to continually train the foundation model with domain corpus as introduced in Section 3.2 Luo et al. (2022). For instance, some models such as BioGPT Luo et al. (2022), ClinicalGPT Wang et al. (2023e), PMC-LLaMA Wu et al. (2023b) and MedAlpaca Han et al. (2023) have been progressively fine-tuned foundation models on the medical related corpus, achieving dramatic medical performance enhancement over the original foundation model. In particular, ChatDoctors Yunxiang et al. (2023) resembles more of a multidisciplinary doctor, built on mixed data instead of solely medical data, capable of conducting patient-doctor dialogues, focusing on comprehensive inquiry services in the real-world scenario. In addition, in the realm of traditional Chinese medicine, HuaTuoGPT2 Chen et al. (2023d) that is fine-tuned on the Baichuan Yang et al. (2023c) with four types of data (distilled instructions/conversations from ChatGPT Ouyang et al. (2022) and real-world instructions/conversations from doctors) and capable of mimicking the diagnostic abilities of doctors provides useful medical information. The experimental results demonstrate the capabilities of HuaTuoGPT2 surpassing those of the GPT-4 Achiam et al. (2023) in the 2023 national medical licensing examination of traditional Chinese medicine. However, the aforementioned models primarily focus on a limited set of high-resource languages and show varying degrees of performance degradation when extended to other languages. Such a phenomenon renders them unreliable, incapable, and insecure for application in linguistically diverse medical environments.
@@ 6.1.1 Medical LLMs in Multilingual Scenarios
To alleviate the issue of multilingualism in the medical domain, prior studies attempt to introduce multilingual medical corpus to enhance the multilingual ability of foundation models as shown in Table 7 and Table 8, Specifically, KBioXLM Geng et al. (2023) adapts XLM-R Conneau et al. (2020b) to the medical domain, which encompasses diverse medical knowledge. However, KBioXLM explores both the training corpus and evaluation data through translation, adapted to only two languages. To further address the limited availability of data beyond English, L2M3 Gangavarapu (2024) integrates Meditron-70B with the Meta SeamlessM4T machine translation system and separately fine-tune two components on four extremely low-resource languages. Garcia et al. Garc´ıa-Ferrero et al. (2024) explore the medical LLM on encoder-decoder architecture based on mT5Xue et al. (2020). In the monolingual setting, both in-context learning (ICL) and SFT demonstrate considerable improvements compared to the foundation models Han et al. (2023); Wu et al. (2023b); Chen et al. (2023c). MMedLM2 Qiu et al. (2024a) underscores medical multilingualism by presenting a 25.5B massive cross-lingual training corpus covering a set of 6 languages. By training with this corpus, the multilingual ability of the model improves in all 6 languages, substantially surpassing the prior models and even rivaling GPT-3.5 and Gemini-1.0-pro. Meanwhile, to further assess the multilingual generalization of medical LLMs, BioMistral Labrak et al. (2024) introduces the first large-scale multilingual medical benchmark of LLMs into 7 languages. Similar to MMedLM2 Qiu et al. (2024a), the Apollo Wang et al. (2024a) introduces a training dataset “Apollo Corpora”, a benchmark “XMedBench” and a collection of models ranging from 0.5B to 7B. To investigate the effectiveness of LLMs in the multilingual medical domain, they build a benchmark “XMedBench”. The comparison on the XMedBench demonstrates that prior multilingual medical LLMs are limited by the PubMed-Central corpus which is constructed based on the translation technique. To alleviate the issue of training data, they propose a training corpus “Apollo Corpora” which is rich with high-quality medical knowledge in each language. The corpus is meticulously collected from the local language and strictly prohibits any form of translation. As a result, the Apollo-7B is fine-tuned on the “Apollo Corpora” and can achieve better performance compared with prior multilingual medical LLMs with 70B parameters.
@@ 6.1.2 Discussion and Challenges
Existing studies achieve remarkable progress in medical LLMs for the multilingual scenario, yet numerous challenges are perverse. First, given the presence of language-specific knowledge that is highly pertinent to the local cultural, historical, political, and regional backgrounds, Wang et al. Wang et al. (2024a) examine whether these language-specific medical knowledge stimulates or deteriorates with each other and investigate whether the model always outperforms its counterpart trained solely on the monolingual language corpus when further trained on the whole multilingual corpus. For instance, the ability of the Spanish model trained in 6 languages surpasses that of a model only trained on a monolingual corpus in Spanish. Despite underlying conflicts between different language-specific medical knowledge and potential biases due to varying data, the performance boost suggests that cross-lingual joint training promotes the performance of medical LLMs, shedding light on the potential of cross-lingual pre-training. Thus, further exploration into the effectiveness of the real-world and pseudo data is yet to be undertaken.

### v2

@@ 6.2 Defense Methods
Only a few studies attempt to address the defense methods in the security LLMs, which can be categorized into open-source and close-source LLMs. Based on the available access to the open-source LLMs, the existing studies [143, 147, 150] enhance the security by fine-tuning the foundation models with security instructions. For closed-source LLMs, previous works [154, 153, 150, 152] defend the risks by auditing the input prompts with various security judgment strategies. However, these simple mechanisms cannot achieve satisfactory performance, as shown in Table 6. The results demonstrate that it is difficult to completely eliminate the generation of unsafe content, regardless of the defense mechanism. And “SmoothLLM” [150] performs better than “Perplexity Filter” [152] but still generates over 10% unsafe content in most scenarios. It indicates that current defense methods for LLMs are in the early stages and require further exploration.
@@ 6.2.1 Defense Methods for Open-Source LLM
Robey et al. [150] propose SmoothLLM, a compatible defense method that acts as a wrapper to smooth the outputs of LLMs by perturbing original attack inputs. Although effective in prohibiting the attacking intent, it also disrupts the semantics of the original input and leads to performance degradation. Deng et al. [143] propose a data augment defense framework SELF-DEFENSE to defend the attacks for both unintentional and intentional scenarios. The key idea is to translate some English seed examples into a target language and then merge the language-specific corpora into the original dataset. Li et al. [147] implement a fine-tuning mitigation method based on their multilingual jailbreak prompts dataset, reducing the attack success rate by 96.2%. Zhou et al. [151] introduce Robust Prompt Optimization (RPO) to LLM security by minimizing the confrontation loss. With less impact on benign cues, the success rate of attacking GPT-4 with GUARD [141] dropped from 96% to 4%.
@@ 6.2.2 Defense Methods for Closed-Source LLM
For closed-source LLM, the defense algorithms mainly focus on dealing with the model inputs. Jain et al. [152] introduce a filter method, judging the danger of the input based on the perplexity. Wu et al. [153] propose a simple and effective defense method SELFDEFEND, which incorporates a shadow stack to check whether harmful prompts exist in the input. Li et al. [154] introduce a Rewindable Auto-regressive INference (RAIN) approach, which has the same evaluation setting as SELFDEFEND but focuses on the output of LLMs. The key component is to leverage another LLM to score the output content and determine whether the output contains harmful content, which achieves good results on LLaMA-base models.
@@ 6.3.1 Jailbreak by Targeting the Multilingual Ability of LLMs
Low-Resource language attack. Although LLMs already have strong multilingual capabilities, they cannot remedy the weakness of low-resource languages in the corpus. The ability of LLMs to identify harmful contents in low-resource languages becomes much worse. Existing studies [145, 143, 142] explore a lot in low-resource languages, while only focusing on GPT-4 with incorporated translation models. Such a pipeline would inevitably introduce noise. Meanwhile, the evaluation standards should be improved to identify the security capacity of LLMs.
The asymmetry between different languages. Language security is highly culturally specific, and each language has its own security vulnerabilities, leading to asymmetry. For example, the word “uso” has insulting connotations only in Basque, thus it needs specific supervised tuning for each language, which is difficult to transfer from other languages. This contradicts the traditional multilingual methods, where knowledge between different languages can be mutually beneficial [264, 265]. Xu et al. [418] empirically validate the presence of multilingual human values within LLMs, encompassing many categories of human values across multiple languages and LLM series. However, the problems of asymmetry between different languages have risen, and no existing studies about it.
Customized encryption and language rules. To bypass the security mechanism of LLMs, an innovative way to create a new language “X” by customizing some language or encryption rules [403, 149, 148] from the original input. This is based on the translation paradigm, which can be viewed as a mapping function to transform character set A to character set B. Thus, exploiting the potential loopholes when aligning LLMs in different languages becomes an important problem.
Misleading information in specific languages. The prior method attempts to assign a word with a specific meaning by embedding a backdoor or directly defining it through dialogue [414, 141]. For instance, Rando et al. [399] insert the word “SUDO” at the end of the query, causing the LLM to misunderstand and bypass the security defenses. It provides insight for attacking the LLMs through multilingual languages. People can introduce words or phrases with specific meanings into queries provided to LLMs. It can lead to the model misunderstanding the meaning of queries, which results in the generation of harmful information. In particular, the polysemous phenomenon that is prevalent in most languages can be used to mislead LLMs into producing prohibited content.
@@ 6.3.2 Robustness of LLMs with Multilingualism
Adversarial training with multilingual samples. Inspired by Mazeika et al. [419], the proposed method enhances defense and robustness using adversarial training. Unlike fine-tuning on fixed datasets with harmful prompts, they propose a Robust Refusal Dynamic Defense to fine-tune the LLM on a dynamic pool of samples, which are continuously updated through a strong optimization-based jailbreak method. It is feasible to incorporate multilingual data into the sample pool using multilingual jailbreak methods which are more practical in attacks, making the data in the sample pool more diverse and advantageous for adversarial training. Besides, adversarial training can be introduced in the pre-training stage or fine-tuning stage of LLMs. As a result, the safety and robustness of LLMs with multilingualism can be enhanced at the root by adversarial training.
Multilingual security alignment with data augmentation. Similar to the approach of enhancing LLMs instruction-following capabilities and human values proposed by SFT and RLHF [1], effective data augmentation is the direct way to improve multilingual alignment. Thus, it is beneficial to explore ways to construct multilingual security datasets facing the localization of security information across languages. Depending on the business or scientific requirements, the datasets can contain different harmful question pairs and languages and then fine-tune the generic LLM with customized security datasets to construct a more robust version of LLMs.
Pre-processing prompts or post-processing outputs. Another approach with fewer resources is to add proxy security efforts in small models outside of LLMs. Before constructing prompts for the LLM, we can consider maintaining a harmful vocabulary list or using another LLM as an evaluator to extract dangerous elements and judge the security of prompts, which can help detect unstable factors. The judgment outputs from LLMs with scores indicate whether the risks are identified in the original inputs and need to be rewritten, similar to RAIN [154]. The general goal is to force LLMs to reject harmful answers or replace them with a rule-based template.
@@ 7 Multi-Domain LLMs in Multilingual Scenarios
The remarkable capabilities of LLMs have facilitated their application across diverse domains, including finance [27, 420, 421, 422, 423, 424], medicine [425, 426, 427, 428, 429, 430, 431], law [432, 433, 434, 435, 436], education [437, 438], transportation [439] etc. These domain-specific LLMs have demonstrated superb capability and promising perspective within associated domains. However, these LLMs are predominantly focused on English, while fewer cater to medium or low-resource languages, which dramatically hinders the utility of LLMs for a global audience. In this chapter, we present the pioneering multilingual studies conducted in the medical and legal domains§§§To the best of our knowledge, at the time we conduct our survey, there are few pertinent works in other domains such as finance and education in multilingual scenarios., and we conclude by offering a comprehensive discussion on the emerging limitations and challenges.
@@
BioMistral [156] 7B Mistral-7B PMC 1.25% $\times$ 3B 10
MMedLM2 [157] 7B InterLM2-7B MMedC 58% $\times$ 25.5B 6
Apollo [158] 0.5/1.8/2/6/7B Qwen/Gemma/Yi Apollo 33.72% $\times$ 2.5B 6
Medical mT5 [160] 0.77/3B mT5 - 66.67% $\times$ 3B 4
L2M3 [159] 70B Meditron& SeamlessM4T - 80% 0.9*5B 5
@@
BioMistral[156] - Multi-choice QA ACC en,ar,zh,fr,de,pt,ru,es
MMedLM2[157] MMedBench Multi-choice QA Rationale eval ACC ROUGE-1/BLEU-1 en,zh,ja,fr,ru,es
Apollo[158] XMedBench $\times$ Multi-choice QA ACC en,zh,fr,es,ar,hi
Medical mT5[160] - $\times$ Seqence labeling Abstractive QA F1 score en, es, it, fr
@@ 7.1 Medical Domain
Previous works have made substantial strides in integrating LLMs into the medical domain [425, 426, 440]. In particular, Med-PaLM2 [441] notably achieves success by passing the US Medical Licensing Examination. A common practice to adapt LLMs to the medical domain is to continually train the foundation model with domain corpus as introduced in Section 3.2 [26]. For instance, some models such as BioGPT [26], ClinicalGPT [442], PMC-LLaMA [443] and MedAlpaca [444] have been progressively fine-tuned foundation models on the medical related corpus, achieving dramatic medical performance enhancement over the original foundation model. In particular, ChatDoctors [445] resembles more of a multidisciplinary doctor, built on mixed data instead of solely medical data, capable of conducting patient-doctor dialogues, focusing on comprehensive inquiry services in the real-world scenario. In addition, in the realm of traditional Chinese medicine, HuaTuoGPT2 [446] that is fine-tuned on the Baichuan [44] with four types of data (distilled instructions/conversations from ChatGPT [1] and real-world instructions/conversations from doctors) and capable of mimicking the diagnostic abilities of doctors provides useful medical information. The experimental results demonstrate the capabilities of HuaTuoGPT2 surpassing those of the GPT-4 [2] in the 2023 national medical licensing examination of traditional Chinese medicine. However, the aforementioned models primarily focus on a limited set of high-resource languages and show varying degrees of performance degradation when extended to other languages. Such a phenomenon renders them unreliable, incapable, and insecure for application in linguistically diverse medical environments.
@@ 7.1.1 Medical LLMs in Multilingual Scenarios
To alleviate the issue of multilingualism in the medical domain, prior studies attempt to introduce multilingual medical corpus to enhance the multilingual ability of foundation models as shown in Table 7 and Table 8, Specifically, KBioXLM [155] adapts XLM-R [447] to the medical domain, which encompasses diverse medical knowledge. However, KBioXLM explores both the training corpus and evaluation data through translation, adapted to only two languages. To further address the limited availability of data beyond English, L2M3 [159] integrates Meditron-70B with the Meta SeamlessM4T machine translation system and separately fine-tune two components on four extremely low-resource languages. Garcia et al. [160] explore the medical LLM on encoder-decoder architecture based on mT5[31]. In the monolingual setting, both in-context learning (ICL) and SFT demonstrate considerable improvements compared to the foundation models [444, 443, 440]. MMedLM2 [157] underscores medical multilingualism by presenting a 25.5B massive cross-lingual training corpus covering a set of 6 languages. By training with this corpus, the multilingual ability of the model improves in all 6 languages, substantially surpassing the prior models and even rivaling GPT-3.5 and Gemini-1.0-pro. Meanwhile, to further assess the multilingual generalization of medical LLMs, BioMistral [156] introduces the first large-scale multilingual medical benchmark of LLMs into 7 languages. Similar to MMedLM2 [157], the Apollo [158] introduces a training dataset “Apollo Corpora”, a benchmark “XMedBench” and a collection of models ranging from 0.5B to 7B. To investigate the effectiveness of LLMs in the multilingual medical domain, they build a benchmark “XMedBench”. The comparison on the XMedBench demonstrates that prior multilingual medical LLMs are limited by the PubMed-Central corpus which is constructed based on the translation technique. To alleviate the issue of training data, they propose a training corpus “Apollo Corpora” which is rich with high-quality medical knowledge in each language. The corpus is meticulously collected from the local language and strictly prohibits any form of translation. As a result, the Apollo-7B is fine-tuned on the “Apollo Corpora” and can achieve better performance compared with prior multilingual medical LLMs with 70B parameters.
@@ 7.1.2 Discussion and Challenges
Existing studies achieve remarkable progress in medical LLMs for the multilingual scenario, yet numerous challenges are perverse. First, given the presence of language-specific knowledge that is highly pertinent to the local cultural, historical, political, and regional backgrounds, Wang et al. [158] examine whether these language-specific medical knowledge stimulates or deteriorates with each other and investigate whether the model always outperforms its counterpart trained solely on the monolingual language corpus when further trained on the whole multilingual corpus. For instance, the ability of the Spanish model trained in 6 languages surpasses that of a model only trained on a monolingual corpus in Spanish. Despite underlying conflicts between different language-specific medical knowledge and potential biases due to varying data, the performance boost suggests that cross-lingual joint training promotes the performance of medical LLMs, shedding light on the potential of cross-lingual pre-training. Thus, further exploration into the effectiveness of the real-world and pseudo data is yet to be undertaken.